OpenStudy (anonymous):

simplify completely 4x2-7x+3/x2+5x-6

OpenStudy (anonymous):

Results: 4x ^4 - 2 x^3 - 6 x^2 + 3 / x^2 the / = Fraction bar not divide.

OpenStudy (callisto):

(4x^2-7x+3)/(x^2+5x-6) = (x-1)(4x-3) / (x+6)(x-1) = (4x-3)/(x+6)